# Long-RVOS: A Comprehensive Benchmark for Long-term Referring Video Object Segmentation

### v1

@@ 5.1 Experiment Setup
Dataset Split. Long-RVOS is a large-scale dataset containing 2,193 videos and 24,689 sentences, which are split into three subsets: a training set of 1,855 videos and 20,722 sentences, a validation set of 113 videos and 1,379 sentences, and a test set of 225 videos and 2,588 sentences.
Evaluation Metrics. We use three kinds of evaluation metrics: the spatial metric $\cal{J}\&\cal{F}$, the temporal metric $\mathrm{tIoU}$ and the spatiotemporal metric $\mathrm{vIoU}$. Long-RVOS provides three types of descriptions: static, temporal and hybrid. We report performance for each type separately and overall. Additionally, we report the FPS for each competitor because efficiency is a major concern for long-video processing.
Implementation Details. We follow the default hyper-parameter settings of ReferDINO [25] and use Swin-Tiny as the backbone. For SAM2 [35], we use the $\mathrm{sam2.1\_hiera\_large}$ version. In MPEG-4 [23], each video clip typically consists of a keyframe and the motion vectors for up to 11 subsequent frames. During training, we randomly sample 6 clips and use 3-frame motion vectors. The input frames are resized to have the longest side of 640 pixels and the shortest side of 360 pixels during training and evaluation. Following the settings on MeViS [7], we do not use referring image segmentation datasets (e.g., RefCOCO/+/g [19, 32]) for pretraining. We train ReferMo on Long-RVOS dataset for 6 epochs, which take 24 hours on 8 Nvidia A6000 GPUs.
@@ 5.2 Benchmark Results
Overall Comparison. We compare ReferMo with six recent RVOS methods on Long-RVOS. All models in comparison are trained on Long-RVOS under consistent experimental settings for fairness. As demonstrated in Table˜2, realistic long-video scenarios remain a significant challenge for current RVOS models. While the SAM2-based methods [26, 5] achieve SOTA performance on existing short-term benchmarks [7, 36, 20], they significantly struggle in Long-RVOS. This suggests that their improvements may primarily stem from SAM2’s superior tracking and segmentation capabilities, rather than better language-guided object understanding. As the videos grow longer and more complex, it becomes more challenging to perform video-language reasoning and distinguish the objects, which leads to their performance degradation. In contrast, our baseline ReferMo integrates the static attributes, short-term dynamics and long-term dependencies to perform object-level visual-language reasoning, achieving significant improvements over existing methods. These findings highlight the need for both frame-level segmentation precision and video-level visual-language understanding to address the long-video challenges in Long-RVOS.
Method Year Static Dynamic Hybrid Overall FPS
@@
SOC [30] 2023 34.8 67.7 28.4 34.9 68.7 28.8 35.1 68.0 28.5 34.9 68.1 28.6 53.8
MUTR [48] 2024 43.0 70.1 36.7 40.2 70.8 34.8 43.2 70.3 37.2 42.2 70.4 36.2 20.4
ReferDINO [25] 2025 50.7 71.9 42.8 45.9 71.9 38.9 49.2 71.5 41.7 48.7 71.7 41.2 46.4
@@
VideoLISA [1] 2024 34.3 69.6 28.9 31.0 69.7 26.9 33.9 69.4 28.6 33.1 69.6 28.2 6.6
GLUS [26] 2025 36.4 68.2 34.3 37.6 68.9 35.8 35.9 68.0 33.9 36.6 68.4 34.6 3.6
SAMWISE [5] 2025 36.6 68.4 29.2 34.3 68.6 28.1 33.8 69.4 28.4 35.6 68.4 28.6 7.0
ReferMo 2025 53.5 71.4 44.0 48.1 71.2 40.1 52.2 71.2 43.6 51.3 71.2 42.6 52.5
Table 2: Comparison of state-of-the-art RVOS models on Long-RVOS test set. FPS is estimated at 360P on Nvidia A6000 GPUs, excluding the video loading time.
Dataset Point Box Mask
MeViS [7] Valid_u 77.3 80.0 80.6
Long-RVOS Valid 53.4 54.5 53.5
Test 52.8 53.9 53.3
(a) Oracle analysis with SAM2.
Model $\mathcal{J}\&\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$
ReferDINO 49.1 47.6 50.6
ReferMo 49.6 48.0 51.2
-w/o motion 47.5 46.0 48.9
(b) Results on the keyframes.
(a) Different motion lengths.
Table 3: Oracle analysis and ablation studies.
Fine-grained Evaluation. Long-RVOS provides three types of text descriptions to enable rigorous evaluation. For most models, the performance for static and hybrid types is comparable and largely better than that for dynamic type. This implies a strong bias in current RVOS models toward static attributes. Across different models, while the $\mathcal{J}\&\mathcal{F}$ scores show significant variance, their $\mathrm{tIoU}$ are relatively consistent. This reveals that existing RVOS models have little performance gap in temporal consistency, highlighting the need for effective tracking mechanisms to handle frequent target disappearance in long-term videos. ReferMo significantly outperforms other models across various types and metrics, except for tIoU, where it is slightly inferior to ReferDINO. We speculate that this is because ReferMo only performs language-guided reasoning on keyframes, resulting in suboptimal object identification on motion frames.
Oracle Analysis. We provide SAM2 with first-frame ground-truth object prompts and evaluate its tracking performance across different datasets. As shown in Table˜3 (a), the oracle results for Long-RVOS (52.8~54.5 $\mathcal{J}\&\mathcal{F}$) are significantly lower than those for MeViS (77.3~80.6 $\mathcal{J}\&\mathcal{F}$). The notable performance gap of nearly 25% demonstrates the long-term challenges in Long-RVOS.
5.3 Ablation Studies
Results on Keyframes. In Table˜3 (b), we compare the performance of ReferMo and ReferDINO [25] on the keyframes. We focus on the spatial metrics since the length of the keyframe sequence is short. Note that ReferDINO is trained on all frames, while our ReferMo is only trained on keyframes. However, ReferMo still outperforms ReferDINO by 0.5% in $\mathcal{J}\&\mathcal{F}$, owing to the integration of motion information. When ablating it, we see a significant 2.1% performance drop in $\mathcal{J}\&\mathcal{F}$. These results encourage further exploration of sparse-frame supervision for RVOS task.
Effect of Motion Information. We investigate the impact of varying the number of motion frames in ReferMo. As shown in Table˜3 (c), the performance without motions is only 49.7 $\mathcal{J}\&\mathcal{F}$. However, even using just one motion frame yields +1.6% $\mathcal{J}\&\mathcal{F}$ improvements. Increasing the motion length to 3 frames improves $\mathcal{J}\&\mathcal{F}$ to 51.3, but further increasing only leads to marginal gains.
6 Conclusion
In this work, we introduce Long-RVOS, a large-scale benchmark for long-term referring video object segmentation, comprising over 2,000 videos averaging 60+ seconds to address the limitations of existing short-term datasets. To enable comprehensive and rigorous evaluation, we provide three types of descriptions and two novel metrics, $\mathrm{tIoU}$ and $\mathrm{vIoU}$. Results on Long-RVOS indicate that current RVOS methods struggle severely in long-video scenarios. Furthermore, we propose ReferMo, a simple motion-enhanced baseline that significantly outperforms existing SOTA methods on long-term videos. We believe that Long-RVOS and ReferMo will provide a foundation for future research to develop robust RVOS models tackling real-world long-form videos.

### v2

@@ 5.1 Experiment Setup
Dataset Split. Long-RVOS is a large-scale dataset containing 2,193 videos and 24,689 descriptions. It is split into three subsets: a training set of 1,855 videos and 20,722 descriptions, a validation set of 112 videos and 1,326 descriptions, and a test set of 226 videos and 2,641 descriptions.
Evaluation Metrics. We use three kinds of evaluation metrics: the spatial metric $\cal{J}\&\cal{F}$, the temporal metric $\mathrm{tIoU}$ and the spatiotemporal metric $\mathrm{vIoU}$. Long-RVOS provides three types of descriptions: Static, Temporal and Hybrid. We report performance for each type separately and overall. We also recommend reporting FPS because efficiency is a major concern for long-video processing.
Implementation Details. We follow the default hyper-parameter settings of ReferDINO [27] and use Swin-Tiny as the backbone. For SAM2 [37], we use the $\mathrm{sam2.1\_hiera\_large}$ version. In MPEG-4 [25], each video clip typically consists of a keyframe and the motion vectors for up to 11 subsequent frames. During training, we randomly sample 6 clips and use 3-frame motion vectors. The input frames are resized to have the longest side of 640 pixels and the shortest side of 360 pixels during both training and evaluation. Following MeViS [9], we do not use image segmentation datasets (e.g., RefCOCO/+/g [21, 34]) for pretraining. We train ReferMo on Long-RVOS dataset for 6 epochs, which take 24 hours on 8 Nvidia A6000 GPUs.
@@ 5.2 Benchmark Results
Overall Comparison. We compare ReferMo with 7 recent RVOS methods on Long-RVOS. For a fair comparison, all the models are trained solely on Long-RVOS, with no external segmentation datasets used. As demonstrated in Table 3, realistic long-video scenarios remain a significant challenge for current RVOS models. While the SAM2-based methods [28, 7, 45] achieve SOTA performance on existing short-term benchmarks, they significantly struggle in Long-RVOS. This suggests that their improvements may primarily stem from SAM2’s superior tracking and segmentation capabilities, rather than better language-object understanding. As videos grow longer and more complex, it becomes more challenging to maintain accurate video-language reasoning and consistently distinguish the objects, which leads to their performance degradation. This issue is more pronounced for those MLLM-based approaches [1, 28, 45], which typically require massive multi-source training data to bridge the gap between reasoning and segmentation. In contrast, our ReferMo demonstrates remarkable data efficiency and inference speed, while achieving significant improvements in long-video understanding.
Table 4: Oracle analysis and ablation studies.
(a) Oracle analysis with SAM2.
Dataset Point Box Mask
MeViS [9] Valid_u 77.3 80.0 80.6
Long-RVOS Valid 54.4 55.9 56.6
Test 54.3 55.6 55.6
(b) Effect of the video decomposition.
Strategy $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}\&\mathcal{F}$
Baseline [27] 48.1 49.7 48.9
+ keyframe. 49.5 50.6 50.0
+ keyframe & motion 50.3 51.8 51.1
(c) Different mask propagation methods.
Method $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}\&\mathcal{F}$
Xmem++ [2] 49.9 51.0 50.4
Cutie [6] 49.6 50.9 50.2
SAM2 [37] 52.2 53.5 52.9
Fine-grained Evaluation. For most models, the highest performance is achieved on the Static type, followed by Hybrid, and the lowest on Dynamic. This implies a strong bias in current RVOS models toward static attributes, as well as significant challenges in dynamic and temporal understanding. Furthermore, while $\mathcal{J}\&\mathcal{F}$ scores vary significantly, $\mathrm{tIoU}$ is relatively stable across methods and types. This indicates that high $\mathcal{J}\&\mathcal{F}$ scores do not necessarily correlate with strong temporal consistency, and the introduction of $\mathrm{tIoU}$ effectively disentangles these aspects. Additionally, the consistently low $\mathrm{vIoU}$ scores across all models suggest that previous evaluations relying solely on frame-averaging metrics may have overestimated the practical robustness of RVOS models. Against this challenging backdrop, our ReferMo showcases consistent performance improvements over SOTA competitors across all types and metrics.
[…]
5.3 Ablation Studies
Effect of Video Decomposition. In contrast to prior RVOS models [27, 52, 32] that directly performs temporal reasoning on the entire video, our ReferMo decomposes the video into clips (keyframe + motion information) to enable local-to-global reasoning. To explore the effect of our strategy, we report performance on the keyframes (before SAM2 tracking) in Table 4(b). The results show that the keyframe-based decomposition strategy surpasses the baseline by 1.1 $\mathcal{J}\&\mathcal{F}$. Further incorporating motion information yields an additional +1.1 $\mathcal{J}\&\mathcal{F}$ gain. Moreover, unlike the baseline, our ReferMo is only trained with keyframe ground truths, yet it achieves much better performance in long-video scenarios. These results encourage further exploration of sparse-frame supervision for RVOS task.
Table 5: Keyframe performance with global interaction.
Global Motion $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}\&\mathcal{F}$
49.8 50.4 49.8
49.5 50.6 50.0
50.0 51.4 50.7
50.3 51.8 51.1
[…]
Effects of Global Interaction and Motion Information. In Table 5, we explore the effect of global interaction. We observe that a naive local-to-global structure only yields a marginal gain of +0.2 $\mathcal{J}\&\mathcal{F}$ (Row 2 vs. Row 1). This is because the sparse keyframes provide insufficient context for global reasoning. In contrast, when we integrate motion features to expand the local window, performance increases significantly by 1.1 $\mathcal{J}\&\mathcal{F}$ (Row 4 vs. Row 2).
Table 6: Overall $\mathcal{J}\&\mathcal{F}$ results at various object occlusion rates.
Method [0, 0.25] [0.25, 0.5) [0.5, 0.75) [0.75, 1]
RGA3 [45] 25.6 17.8 19.1 10.4
MUTR [52] 47.4 38.5 30.8 17.4
ReferDINO [27] 53.3 45.0 36.7 25.6
SAMWISE [7] 39.9 39.3 38.8 38.0
ReferMo (Ours) 54.6 50.6 46.5 39.7
Robustness of Keyframe Methods. As a keyframe-based approach, ReferMo may encounter challenges when target objects are absent from selected keyframes. To evaluate its robustness, we present the results under varying object occlusion rates in Table 6. The results show that ReferMo consistently outperforms all competitors across all occlusion brackets. Moreover, as the occlusion rate increases, ReferMo maintains a consistent performance advantage over most methods. Although its leading margin over SAMWISE [7], which uses a streaming post-correction mechanism, narrows at high-occlusion scenarios, ReferMo’s overall performance is significantly superior. Therefore, despite relying soly on keyframe reasoning, ReferMo remains sufficiently robust in most non-extreme cases.
6 Conclusion
In this work, we introduce Long-RVOS, a large-scale benchmark for long-term referring video object segmentation, comprising over 2,000 videos averaging 60+ seconds to address the limitations of existing short-term datasets. To enable comprehensive and rigorous evaluation, we provide three types of descriptions and two novel metrics, $\mathrm{tIoU}$ and $\mathrm{vIoU}$. Results on Long-RVOS indicate that current RVOS methods struggle severely in long-video scenarios. Furthermore, we propose ReferMo, a simple motion-enhanced baseline that significantly outperforms existing SOTA methods on long-term videos. We believe that Long-RVOS and ReferMo will provide a foundation for future research to develop robust models tackling real-world long videos.
Supplementary Material
Table 7: Definitions of the video attributes.
   Attribute    Full Name    Definition
   POC    Partial Occlusion    The target object is partially occluded in the sequence.
   FOC    Full Occlusion    The target object is fully occluded in the sequence.
   OV    Out-of-view    The target leaves the video frame completely.
   LRA    Long-term Reappearance    Target object reappears after disappearing for at least 100 frames.
   VC    View Change    Viewpoint affects target appearance significantly.
   ARC    Aspect Ratio Change    The ratio of bounding box aspect ratio is outside the range [0.5, 2].
   SV    Scale Variation    The ratio of any pair of bounding-box is outside of range [0.5,2.0].
   CM    Camera Motion    Abrupt motion of the camera.
   MB    Motion Blur    The boundary of target object is blurred because of camera or object fast motion.
Table 8: The percentage (%) of videos featuring specific attributes.
   Dataset    POC    FOC    OV    LRA    VC    ARC    SV    CM    MB
   MeViS [9]    54.8    15.1    28.7    0.1    10.0    88.2    78.7    49.2    18.8
   Long-RVOS (Ours)    60.5    36.2    61.0    11.5    25.9    96.2    93.6    60.7    28.7
7 More Dataset Statistics
To further highlight the challenges posed by Long-RVOS, we present a statistical analysis of video attributes, with definitions provided in Long-RVOS: A Comprehensive Benchmark for Long-term Referring Video Object Segmentation. As shown in Long-RVOS: A Comprehensive Benchmark for Long-term Referring Video Object Segmentation, compared to the current largest dataset MeViS [9], Long-RVOS involves numerous long-video challenges, including frequent object occlusion (POC, FOC, and OV) and long-term object disappearance-reappearance (LRA). In addition, the videos in Long-RVOS exhibit significant object motion (CM and MB) and appearance changes (VC, ARC and SV), making the dataset more akin to real-world scenarios.
8 More Implementation Details
Motion Extraction. Following Video-LaVIT [20], we rely on motion vectors instead of the expensive dense optical flow. Formally, given a video clip, we partition each frame into $16\times 16$ non-overlapping macroblocks. Then, motion vectors $\vec{m}$ of the $t$-th frame are estimated by matching the macroblocks between the adjacent frames $I_{t}$ and $I_{t-1}$:
$\vec{m}(p,q)=\arg\min_{i,j}\|I_{t}(p,q)-I_{t-1}(p-i,q-j)\|,$ (6)
where $I(p,q)$ denotes the pixel values of the macroblock at location $(p,q)$, and $(i,j)$ denotes the coordinate offset between the centers of the two macroblocks. Empirically, the extraction of motion vectors runs at 748 FPS on our device (Intel(R) Xeon(R) Gold 6226R CPU @ 2.90GHz), enabling real-time processing of long videos.
Global Interaction. This module performs temporal attention over the inter-frame object features, enabling temporal reasoning and understanding. Since this is a common module in RVOS approaches [32, 9, 27], we follow the object-consistent temporal enhancer (OTE) of ReferDINO [27] rather than designing a new one from scratch. For clarity, we briefly revist OTE here. Given $T$-frame object features $\{O_{t}\}_{t=1}^{T}$ where $O_{t}\in\mathbb{R}^{N_{q}\times d}$, OTE utilizes the Hungarian algorithm [24] to align the $N_{q}$ objects between adjacent frames based on the pairwise cosine similarity. After that, it performs temporal self-attention over the aligned object features and cross-attention with the sentence features $\widetilde{E}$. We refer the readers to the original paper [27] for additional details.
Training. Unlike previous RVOS methods, ReferMo relies only on the keyframe ground-truth annotations for efficient training. Formally, given a text description and a video of $T_{c}$ clips, ReferMo outputs the prediction sequences $\{\mathbf{p}_{i}\}_{i=1}^{N_{q}}$ for the $N_{q}$ object queries, where each sequence $\boldsymbol{p}_{i}=\{\hat{\boldsymbol{s}}^{t}_{i},\hat{\boldsymbol{b}}^{t}_{i},\hat{\boldsymbol{m}}^{t}_{i}\}_{t=1}^{T_{c}}$ describes the binary classification probability, bounding box and mask of the $i$-th object query on $t$-th keyframe. Our training pipeline follows the practice in previous approaches [48, 32, 27]. Suppose $\boldsymbol{y}=\{\boldsymbol{s}^{t},\boldsymbol{b}^{t},\boldsymbol{m}^{t}\}_{t=1}^{T_{c}}$ as the ground truth of keyframes, we select the prediction sequence with the lowest matching cost as the positive and assign the remaining sequences as negative. The matching cost is defined as follows:
$\begin{split}\mathcal{L}_{\text{total}}\left(\boldsymbol{y},\boldsymbol{p}_{i}\right)=&\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}\left(\boldsymbol{y},\boldsymbol{p}_{i}\right)+\lambda_{\text{box}}\mathcal{L}_{\text{box}}\left(\boldsymbol{y},\boldsymbol{p}_{i}\right)\\ &+\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}\left(\boldsymbol{y},\boldsymbol{p}_{i}\right).\end{split}$ (7)
The matching cost is computed on individual frames and normalized by $T_{c}$. Here, $\mathcal{L}_{\text{cls}}$ is the focal loss that supervises the binary classification prediction. $\mathcal{L}_{\text{box}}$ sums up the L1 loss and GIoU loss. $\mathcal{L}_{\text{mask}}$ is the combination of DICE loss, binary mask focal loss and projection loss [44]. $\lambda_{\text{cls}}$, $\lambda_{\text{box}}$ and $\lambda_{\text{mask}}$ are scalar weights of individual losses. The model is optimized end-to-end by minimizing the total loss $\mathcal{L}_{\text{total}}$ for positive sequences and only the classification loss $\mathcal{L}_{\text{cls}}$ for negative sequences.
Inference. ReferMo produces instance mask for the referring object on keyframes and then employs SAM2 [37] for subsequent frames. Specifically, for the prediction sequences $\{\mathbf{p}_{i}\}_{i=1}^{N_{q}}$, we select the best sequence with the highest average classification score as follows:
$\sigma=\underset{i\in[1,N_{q}]}{\arg\max}\frac{1}{T_{c}}\sum\nolimits_{t=1}^{T_{c}}\hat{\boldsymbol{s}}^{t}_{i}$ (8)
Then, the output mask sequence on keyframes is formed as $\{\boldsymbol{m}_{\sigma}^{t}\}_{t=1}^{T_{c}}$. For the $t$-th video clip, we use the keyframe prediction $\boldsymbol{m}_{\sigma}^{t}$ as the mask prompt for SAM2, which tracks the target across the subsequent frames within the clip.
9 Validation Results
In Table 9, we present the benchmark results on Long-RVOS validation set. The results show that our ReferMo achieves consistent improvements over previous RVOS methods, especially those built on SAM or SAM2.
Table 9: Comparison on Long-RVOS valid set. FPS is estimated at 360P on Nvidia A6000 GPUs, excluding the video loading time.
Method Static Dynamic Hybrid Overall FPS
@@ 9 Validation Results
SOC [32] NeurIPS’23 38.7 73.1 34.9 37.8 74.6 34.1 37.8 74.3 34.5 38.1 74.0 34.5 53.8
MUTR [52] AAAI’24 44.1 73.5 40.3 42.0 75.2 38.9 43.5 74.6 40.2 43.2 74.4 39.8 20.4
ReferDINO [27] ICCV’25 52.5 74.2 48.2 46.7 75.2 42.9 49.3 74.8 45.4 49.6 74.7 45.6 46.4
@@
VideoLISA [1] NeurIPS’24 17.3 66.8 12.7 12.9 72.6 6.8 12.1 72.3 6.0 14.1 70.5 8.6 6.6
GLUS [28] CVPR’25 24.4 62.8 20.8 26.1 64.7 23.1 24.1 63.5 20.6 24.8 63.7 21.5 3.6
SAMWISE [7] CVPR’25 42.3 61.5 31.2 40.7 63.3 31.4 40.6 65.8 31.2 41.2 63.5 31.2 7.0
RGA3 [45] ICCV’25 21.1 61.0 15.4 22.3 62.8 17.5 21.1 61.8 16.4 21.5 61.8 16.4 8.7
ReferMo (Ours) 56.7 74.0 49.4 50.7 74.2 43.4 53.7 74.7 47.4 53.7 74.3 46.8 52.5
10 More Ablation Studies
Effectiveness of Gating Image-Motion Fusion. ReferMo employs the spatial-aware gating (SG) and channel-aware gating (CG) mechanisms in image-motion fusion to avoid undesired motion noise. As shown in Table 10, directly concatenating keyframe and motion features leads to a performance collapse. This is because RVOS requires per-frame fine-grain perception, while directly integrating motion features can introduce significant object-irrelevant noise. By applying these two gating strategies, ReferMo effectively alleviates such noise while preserving only the motion cues that highlight target objects, thereby yielding significant performance gains.
Table 10: Keyframe results of different image-motion fusion approaches.
SG CG $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}\&\mathcal{F}$
28.8 28.7 28.7
46.9 46.3 46.6
49.5 50.4 50.0
50.3 51.8 51.1
Table 11: Overall $\mathcal{J}\&\mathcal{F}$ results for different description lengths.
Method <<10 [10, 20] >>20
RGA3 [45] 23.8 22.6 19.8
MUTR [52] 42.5 43.1 38.2
SAMWISE [7] 40.1 41.5 40.8
ReferDINO [27] 49.2 49.2 44.3
ReferMo (ours) 53.6 53.6 48.5
Effect of Description Length. We evaluate the impact of varying description lengths and present the results in Table R1. As description length increases, slight performance declines are observed across models. However, our ReferMo consistently outperforms existing methods across different description lengths.
Table 12: Overall $\mathcal{J}\&\mathcal{F}$ results by event complexity.
Method Single-event Two-event Multi-event
RGA3 [45] 23.0 22.6 19.2
MUTR [52] 42.7 40.0 36.0
SAMWISE [7] 40.6 41.7 38.0
ReferDINO [27] 48.4 44.7 37.2
ReferMo (ours) 52.9 48.0 40.5
Effect of Multi-event Videos. To explore the impact of event number in a video on model performance, we categorized the samples into single-event, two-event, and multi-event groups based on the keywords (e.g., then, finally, ultimately) in descriptions. As shown in Table 12, performance across models declines as the event number increases, yet our ReferMo consistently outperforms existing methods. Also, these results highlight the significance of our long-term benchmark for evaluating the capabilities of RVOS models in understanding complex event sequences.
Figure 8: Qualitative comparison of our ReferMo with the SOTA model ReferDINO [27]. ReferMo performs better in long-term object consistency and segmentation quality.
11 Visualization
In Figure 8, we provide the qualitative comparisons with the SOTA model ReferDINO [27] on Long-RVOS. These examples involve multiple long-term challenges, such as object occlusion, disappearance-reappearance and view changes. The results clearly show the effectiveness of our baseline ReferMo in long-term object consistency and segmentation quality.
12 Limitations and Future Work
In this work, we chose to begin with description-based RVOS because it is commonly used in current video applications and this task remains far from being solved. It is promising to broaden the benchmark scope to support more tasks, such as reasoning RVOS [1, 51], semi-supervised VOS [36, 50, 10], interactive VOS [23, 37] and audio-guided VOS [52]. Besides, while our benchmark covers a variety of objects, it currently does not include background stuff classes (e.g., sky and river), which could be incorporated in future work for covering more diverse scenarios.